OpenStudy (anonymous):

Does anyone have some practice problems for using functions within functions, like f(g(x))??

OpenStudy (anonymous):

yes what is your question

OpenStudy (anonymous):

like just some problems?

OpenStudy (anonymous):

sure ask away :-)

OpenStudy (anonymous):

okay let me just make one up f(x)= 3x G(x)= x + 4 Is f(g(x))= 3x+12 and g(f(x)) 3x+7?

OpenStudy (anonymous):

Sure, here is a practice problem. What is g(f(x)) if f(x)=3x-5, g(x)=13. Tell me if you need more practice problems

OpenStudy (unklerhaukus):

@amcdowns \[f(x)= 3x\qquad\qquad g(x)= x + 4\] \[ f(g(x)) = 3(x+4) =3x+12 \checkmark\]but \[g(f(x)) 3x+7=(3x)+4=3x+4\checkmark\quad\neq3x+7\]?

OpenStudy (anonymous):

The f(g(x)) is correct, but the g(f(x)) is not correct, you would just have the answer to be 3x+4, there doesn't need to be a 7 in there.

OpenStudy (anonymous):

Okay thanks! So let me make another: F(x)= 2x+8 g(x)= x+4 f(g(x))= 2x+16 g(f(x))= 2x+12 Are these right??

OpenStudy (anonymous):